OpenStudy (anonymous):

I need help on evaluate expressions 6th grade work

OpenStudy (texaschic101):

what ya got ?

OpenStudy (texaschic101):

hello ??

OpenStudy (anonymous):

yea

OpenStudy (anonymous):

oh ive got a problem 13-2X

OpenStudy (texaschic101):

did it give you a value for x ?

OpenStudy (anonymous):

um x= -5 -1 0 and 3

OpenStudy (texaschic101):

13 - 2(-5) = 13 + 10 = 23 13 - 2(-1) = 13 + 2 = 15 13 - 2(0) = 13 - 0 = 13 13 - 2(3) = 13 - 6 = 7

OpenStudy (texaschic101):

just sub in your numbers for x. and solve

OpenStudy (texaschic101):

is that all you need ?

OpenStudy (anonymous):

no I need more help

OpenStudy (anonymous):

6x divided by -3 and x is the same as the other problem

OpenStudy (texaschic101):

6x/-3 (6(-5)) / -3 = -15/-3 = 5 (6(-1)) / -3 = -6/-3 = 2 (6(0)) / -3 = 0/-3 = 0 (6(3)) / -3 = 18/-3 = -6

OpenStudy (texaschic101):

wait...messed up on the first one... (6(-5)) / -3 = -30/-3 = 10

OpenStudy (anonymous):

in need help on another problem

OpenStudy (texaschic101):

is it the problem you messaged me with ?